OpenStudy (anonymous):

partial fraction of dv/v(v+2)=? please make step by step solution.waiting

OpenStudy (anonymous):

\[ \frac{1}{v(v+2)}=\frac Av+\frac B{v+2} \]

OpenStudy (anonymous):

\[ 1 = A(v+1)+B(v) = (A+B)v+A \]

OpenStudy (anonymous):

\[ (0)v + 1 = (A+B)v + A\\ \implies \\ 0=A+B\\ 1=A \]

OpenStudy (anonymous):

in question is 2 not 1

OpenStudy (anonymous):

thanks